Circular No. 8780

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2006 S12, C/2006 S13, C/2006 T2–T9 (SOHO) Additional apparent comets have been found on SOHO website images

(cf. IAUC 8774), all being Kreutz sungrazers except C/2006 T6 (Meyer group). K. Battams notes that C/2006 S12 was small, stellar in appearance, and around mag 6 in C3 coronagraph images, while it appeared slightly diffuse with no tail in C2 images. C/2006 S13 was very small and around mag 7 in C3 frames, and it appeared small, quite diffuse, and extremely faint in C2 exposures. C/2006 T2 was small, stellar in appearance, and around mag 6 in C3 images, and quite diffuse and faint in C2 frames. C/2006 T3 was tiny, stellar in appearance, and of mag ∼ 7.5 in C3 exposures, but it appeared very diffuse and faint in C2 images. C/2006 T4 (which was also found by T. Chen) was diffuse and extremely faint (mag 7.5–8.0) in C3 images, and extremely diffuse and faint in C2 frames. C/2006 T5 was slightly diffuse, extremely small, and faint (mag ∼ 8). C/2006 T6 was tiny and stellar in appearance in all exposures, being very faint (mag 7.5–8) in C3 frames and fading rapidly in C2 images. C/2006 T7 and C/2006 T9 were extremely faint (mag ∼ 8.5) and diffuse with no tail. C/2006 T8 was tiny, stellar in appearance, and of mag ∼ 7.5 in C3 images, while C2 frames showed it to be faint and diffuse with a hint of tail.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 S12 Sept. 26.763 11^{h}48.^{m}2 − 2^{o}50′ C3/2 RM 2006-W64
C/2006 S13  30.433 12 11.9 − 3 48 C3/2 HS 2006-W64
C/2006 T2 Oct. 2.113 12 11.4 − 4 52 C3/2 RM 2006-W64
C/2006 T3  3.263 12 21.6 − 4 59 C3/2 HS 2006-W64
C/2006 T4  3.279 12 21.5 − 5 01 C3/2 GS 2006-W64
C/2006 T5 6.314 12 40.7 − 5 43 C2 WX 2006-W65
C/2006 T6  9.904 12 49.9 − 5 28 C3/2 BZ 2006-W65
C/2006 T7 10.379 12 54.7 − 7 18 C2 BZ 2006-W65
C/2006 T8  15.429 13 08.2 − 9 54 C3/2 TH 2006-W65
C/2006 T9  15.881 13 15.6 − 9 35 C2  TH 2006-W65

COMET C/2006 W3 (CHRISTENSEN) Improved orbital elements (cf. IAUC 8777) from MPEC 2006-W76:

T = 2009 Apr. 23.508 TT  ω = 142.^{o}308
Ω = 112.317 }2000.0
q = 2.38612 AU  i = 125.410
2006 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green